OpenStudy (anonymous):

Is there a closed form formula that yields the number of digits in a decimally-based number ? A special function ?

OpenStudy (anonymous):

3sec^4 x - 10sec^2 x - 8 = 0, what is sin x

OpenStudy (anonymous):

Don't understand the question. I give u 4.35, u want a formula that predicts that there are 3 digits in 4.35..

OpenStudy (anonymous):

If you want to know how many digits are in the integer x: \[\lfloor \ln(x)/\ln(10) \rfloor + 1\]

OpenStudy (anonymous):

Or of course, to compensate for negative values of x: \[\lfloor \ln(\left| x \right|)/\ln(10) \rfloor + 1\]

OpenStudy (anonymous):

The only problem is it doesn't give an answer for x = 0.

OpenStudy (anonymous):

Many thanks. (If this work for strictly positive integer, I'm OK)

OpenStudy (anonymous):

Any time.